### v1

# Comparative Analysis of Weather-Based Indexes and the Actuaries Climate IndexTM for Crop Yield Prediction

Abstract
Climate change poses significant challenges to agriculture sector, affecting both crop productivity and economic stability. This study investigates the utility of climate indexes in predicting crop yields under diverse climatic conditions. Using generalized statistical models and machine learning algorithms, this research compares the predictive performance of long-established weather-based indexes with the Actuaries Climate IndexTM (ACI), a newer entrant in the field, showcasing its potential for real-time applications and broader applicability across various sectors. To enhance model reliability and address multicollinearity among weather-related variables, the study incorporates both principal component analysis and functional principal component analysis. A total of 22 models, each constructed with different explanatory variables, highlight the significant impact of wind speed and sea-level changes alongside temperature and precipitation on crop yield variability. Furthermore, the ACI exhibits strong predictive capabilities, offering valuable insights for climate risk assessment and decision-making in agriculture and insurance-related sectors.
Keywords actuaries climate index  $\cdot$ principal component analysis  $\cdot$ crop yield  $\cdot$ weather-based indexes  $\cdot$ machine learning
1 Introduction
Climate change refers to long-term shifts in temperature and weather patterns, primarily driven by human-induced greenhouse gas emissions. Recent reports from the Intergovernmental Panel on Climate Change highlight the significant impact of human activities on global temperatures over the past few decades (IPCC 2022; Masson-Delmotte et al. 2021). One of the most concerning consequences of climate change is the increase in the frequency and intensity of heatwaves, which exacerbate heat-related illnesses and deaths. WHO (2014) projects that climate change could result in an additional 250,000 deaths annually between 2030 and 2050. Historical events underline the severity of this issue. For instance, in August 2003, France endured its hottest summer in 50 years, leading to over 14,000 fatalities (Kovats et al. 2004). Similarly, a 2010 heatwave in Quebec, Canada, caused a 33% spike in daily mortality rates and hospital emergency visits (Bustinza et al. 2013).
The financial risks for the insurance industry is heightened by climate change, primarily due to the increasing frequency and severity of extreme weather events such as hurricanes, floods, and wildfires. Gall et al. (2011) highlight that direct losses from natural hazards in the United States (U.S.) are steadily increasing and could reach \$300 to \$400 billion within a single decade. A specific example is the California wildfires of 2018, which Wang et al. (2021) estimated to have caused \$148.5 billion in damages, including capital losses, health costs, and indirect economic disruptions. More recently, the wildfires in the Los Angeles area have the potential to inflict up to \$250 billion in economic losses, making them one of the most devastating natural disasters in U.S. history (Danielle 2025). From a broader perspective, Zhou et al. (2024) discuss the profound impact of climate change across various insurance sectors, including property/casualty, health, life, and crop insurance. They underscore the urgency for standardized methodologies to measure and quantify climate risks, ensuring that the industry can adapt to the sustainability challenges posed by a rapidly changing climate.
Climate change also leads to challenges like reduced crop yields and water scarcity, posing significant threats to global food security. Research by Lobell and Gourdji (2012) predicts that global warming could lower global crop yields by approximately 1.5% per decade, with potential reductions reaching as high as 4%. Several other studies reach similar results. For instance, Hu et al. (2024) observed that a 1°C rise in global temperatures results in a 6-8% average decrease in yields of major crops worldwide. Cai et al. (2009) project a 23-24% reduction in corn yields in central Illinois, U.S., by 2055, largely driven by increased water stress. Similarly, Adhikari et al. (2015) examine the impact of climate change on crops in Eastern Africa, concluding that the yields of major grains are significantly affected negatively, with wheat being particularly vulnerable—up to 72% of current yields are projected to decline.
Climate impact is mostly measured and interpreted by indexes. In the literature, various indexes such as the Climate Extremes Index (Karl et al. 1996), Greenhouse Climate Response Index (Karl et al. 1996), Palmer Drought Severity Index (Palmer 1965), Cooling Degree Days (Jewson and Brix 2005), Heating Degree Days (Jewson and Brix 2005) and Cumulative Rainfall Index (Odening et al. 2007) have been developed to quantify climate change, enabling researchers and policymakers to assess systematically its impact. These indexes often focus on critical variables such as temperature extremes, amounts of precipitation, drought severity, summarizing them into measurable indicators that facilitate the analysis of climate patterns and their consequences.
In recent years, the Actuaries Climate IndexTM (ACI) has emerged as a pivotal tool for understanding and monitoring climate risks, offering valuable insights for insurance companies, governments, and the general public. Developed in 2016 by the four major North American actuarial societies, the ACI (2018) uses historical meteorological data from the United States and Canada to track extreme conditions in air temperature, precipitation, drought, wind, and sea-level rise. Its practical methodology has quickly been adapted for use in other countries. Curry (2015) demonstrates that the ACI methodology is applicable to regions like the United Kingdom and Europe, provided appropriate and sufficient meteorological data are available. The Australian Actuaries Climate Index (AACI 2018), which is tailored to Australian meteorological conditions, was introduced in 2018. The Turkish Actuaries Climate Index was developed in 2022 by Nevruz et al. (2022), and focuses on climate patterns specific to the Ankara region in Türkiye. Similarly, Zhou et al. (2023) launched the Iberian Actuarial Climate Index in 2023, marking the first application of the ACI framework to the European continent. It encompasses Spain and Portugal, offering localized insights into climate variability. Most recently, Garrido et al. (2023) adapt the ACI methodology to French data, creating the French Actuarial Climate Index. The continued adoption of the methodology worldwide demonstrates a growing global interest in effectively measuring and understanding climate risks.
This study aims to compare the Actuaries Climate IndexTM, with long-established weather-based indexes frequently used in energy and derivatives markets, by using regression models: generalized linear model (GLM) and generalized additive model (GAM), as well as machine learning algorithms: extreme gradient boosting (XGB) and light gradient boosting machine (LGBM). By evaluating these indexes based on their performance in predicting crop yields, a critical factor in the agricultural sector, the study demonstrates the robustness of the ACI, showcasing its potential for real-time applications and broader applicability across various sectors. Existing literature demonstrates the applicability of the ACI in various fields, particularly insurance and finance. Jiang and Weng (2019) analyze ACI’s predictive power for stock returns of agriculture-related companies in the U.S. and Canada, concluding that ACI could support a profitable trading strategy. Pan et al. (2022) assess ACI’s effectiveness in predicting corn yields in the Midwest region of the U.S. using linear and probit regression models, finding satisfactory predictive power. Zhou and Vilar-Zanón (2024) explore the relationship between the Spanish Actuarial Climate Index (SACI) and hailstorm-related losses in the wine grape insurance industry through linear and quantile regression models, noting a potential rise in monthly losses corresponding to future SACI increases.
Beyond comparing the predictive power of the indexes, this paper investigates the individual contribution of each weather index to model performance by fitting 22 distinct models, each constructed with different explanatory variables. This structured approach allows for a detailed assessment of how the inclusion of diverse weather indexes enhances the models’ predictive accuracy and its explanatory power. Since all explanatory variables are weather-based, they may share common weather driven patterns, potentially leading to multicollinearity issues in regression models. Additionally, achieving comparable model performance with fewer explanatory variables is both practical and desirable. To address these challenges, dimension reduction techniques specifically, principal component analysis (PCA) and functional principal component analysis (FPCA) are applied to identify the most influential variables in the data sets. This approach not only mitigates multicollinearity issues but also simplifies the models, improving efficiency and interpretability.
The outcomes indicate that both ACI components and weather-based indexes exhibit robust predictive capabilities for crop yields (corn, wheat, soybeans) in the U.S. market. Furthermore, the findings reveal that incorporating indexes representing additional weather conditions—such as drought, wind, and sea-level—alongside traditional factors like air temperature and precipitation, substantially enhances model performance. This highlights the importance of a comprehensive approach that considers a broader spectrum of climatic variables to improve the accuracy and reliability of crop yield predictions.
The paper is organized as follows: Section 2 summarizes the climate indexes used in the study. Section 3 describes the proposed two-stage methodology and presents the framework of the 22 models developed for crop yield prediction. The properties of the data and the results of the analyses are provided in Section 4. Finally, Section 5 concludes the paper with a summary of the key findings and their implications. Additionally, regression models, machine learning algorithms and dimensionality reduction techniques used in the analyses are reviewed in the Appendix A.
2 Climate Indexes
We focus on weather-based indexes (WBI) such as Cooling Degree Day (CDD), Heating Degree Day (HDD) and Cumulative Rainfall Index (PRE), which are directly tied to temperature and rainfall. These indexes are particularly advantageous due to their reliance on single weather variables, making them simple to understand, straightforward to compute and convenient for comparative analysis with the Actuaries Climate IndexTM. Detailed formulations of these weather-based indexes and outlines off the corresponding components within the ACI framework are given below.
2.1 Weather-Based Indexes
CDD and HDD are temperature-based indexes, with CDD capturing warm weather conditions and HDD representing cool weather conditions. PRE, on the other hand, quantifies rainfall amounts. These indexes are initially calculated for daily time periods, and their values for longer intervals, such as monthly or meteorological seasons, are derived by summing the daily data. Briefly, CDD measures how much the outside air temperature exceeds a base level temperature (usually taken as 65°F or 18°C), and HDD measures the opposite, the temperature falls below a base level (Jewson and Brix 2005).
Let $T_{i}^{\text{max}}$ and $T_{i}^{\text{min}}$ denote the maximum and minimum temperatures measured on day $i$. The average temperature for day $i$ is defined as,
$T_{i}=\frac{T_{i}^{\text{max}}+T_{i}^{\text{min}}}{2}.$ (1)
If the average temperature is higher than 65°F for CDD, and lower than 65°F for HDD, the difference becomes the $CDD_{i}$ and $HDD_{i}$ value for that day, respectively, given by,
$CDD_{i}=max\{T_{i}-65,0\},$ (2)
$HDD_{i}=max\{65-T_{i},0\}.$ (3)
On the other hand, PRE represents the total volume of rainfall accumulated over a specified time period (Odening et al. 2007), and is calculated by summing all daily rainfall measurements, expressed in millimeters, for the chosen time frame, such as a month or a meteorological season, described as
$PRE=\sum_{t=1}^{N}r_{t},$ (4)
where $r_{t}$ denotes the daily rainfall amount on day $t$, and $N$ is the total number of days in the specified period.
2.2 Actuaries Climate IndexTM
The ACI is designed to improve understanding of climate trends and their potential impact, incorporating various climate variables that signify extreme weather conditions. It comprises six components: high temperature (T90), low temperature (T10), precipitation (P), drought (D), wind speed (W), and sea-level (S). Since each component reflects different weather conditions, the ACI uses standardized values for all components to integrate them into a single index. Standardization is based on the mean and standard deviation of each component during the designated reference period of 1961 to 1990 (ACI 2018).
The components of the ACI used in our comparison analyses are briefly defined as follows:
  1. (i)
    T90 represents the percentage of days in a month when daily temperatures exceed the 90th percentile of the corresponding days in the reference period.
  2. (ii)
    T10 reflects the percentage of days when daily temperatures fall below the 10th percentile of the reference period.
  3. (iii)
    P, denoted as Rx5day, measures the maximum amount of rainfall recorded over any five consecutive days within a month.
The standardized values of each component are computed both on a monthly and seasonal basis. As described in Zhou et al. (2023), for each month ($j$) and year ($k$), these are:
$T90_{STD}(j,k)=\frac{T90(j,k)-\mu_{T90(j)}}{\sigma_{T90(j)}},$ (5)
$T10_{STD}(j,k)=\frac{T10(j,k)-\mu_{T10(j)}}{\sigma_{T10(j)}},$ (6)
$P_{STD}(j,k)=\frac{Rx5day(j,k)-\mu_{Rx5day(j)}}{\sigma_{Rx5day(j)}}.$ (7)
Here, $\mu_{\cdot(\cdot)}$ and $\sigma_{\cdot(\cdot)}$ are the mean and standard deviation of corresponding components for the reference period, respectively. The composite ACI is derived by averaging the standardized values of all six components, with the T10 component being assigned a negative value to account for its inverse relationship with warm weather conditions, and expressed as:
$ACI=\frac{T90_{STD}-T10_{STD}+P_{STD}+D_{STD}+W_{STD}+S_{STD}}{6}.$ (8)
Detailed information of ACI can be found in ACI (2018).
3 Methodology
The primary objective of this study is to assess the robustness of the ACI as a climate impact measure in the agriculture sector. Given that ACI is a relatively recent addition to the literature, its performance in predicting crop yields is compared with the performance of long-established weather-based indexes. By evaluating the predictive capabilities of climate indexes, the study seeks to demonstrate ACI’s potential for real-time applications and its broader utility across various sectors.
The analyses are conducted in two stages. The first stage seeks similarities between WBI and ACI components. To do so, we first check how strongly correlated these two indexes are and define their empirical cumulative distribution functions. The assessment is conducted across three groups, each representing different weather conditions, enabling a systematic and detailed comparison of the two types of indexes, WBI and ACI. The second stage focuses on testing and comparing the predictive performance of climate indexes on crop yields. Crop yields are predicted by using regression models and machine learning algorithms that comprising a total of 22 distinct models, each constructed with different explanatory variables. This approach allows for a detailed analysis of the predictive performance of WBI and ACI, ensuring that variations in their predictive capabilities are systematically evaluated.
3.1 First Stage
Climate indexes that measure similar weather conditions are grouped into categories to facilitate meaningful comparisons. The grouping structure is based on the specific meteorological conditions represented by each index, such as warm weather, cool weather, and precipitation. Despite differences in their calculation methods, CDD defined in equation 2 and T90 in 5 measure warm weather conditions, HDD in 3 and T10 in 6 capture cool weather conditions, and PRE in 4 and P in 7 represent precipitation severity. This grouping ensures a systematic and logical comparison between the indexes, and its structure is summarized in Table 1.
Table 1: Structure of groups
Group
Weather-Based
Indexes (WBI)
Components
of the ACI
Meteorological
Season
Measurement
1 CDD T90 Summer Warm weather
2 HDD T10 Winter Cool weather
3 PRE P Spring Precipitation
For the comparisons, meteorological seasonal values (three months ending February, May, August, November) are used instead of monthly values. This approach aligns with the practical applications of climate indexes. For instance, CDD is commonly used to estimate energy demand during summer months, particularly for cooling, while HDD is used to evaluate heating needs during winter months (Spinoni et al. 2018). Regarding precipitation, its timing plays a critical role in agricultural productivity. Wang et al. (2016) highlight that wheat yields are especially sensitive to rainfall during the rainy season, while Peltonen-Sainio et al. (2011) point out that precipitation before harvest often negatively impacts crop yields. Therefore, in the precipitation group, index comparisons are performed for the spring months to encompass the rainy season preceding the general harvest period, which typically occurs during the summer months for most crops.
The dependence between the indexes are analyzed using the Spearman rank correlation coefficient (Zar 2005), a non-parametric measure that evaluates the strength and direction of the monotonic relationship between two variables. Distribution comparisons are performed using the Kolmogorov-Smirnov (K-S) test (Massey Jr 1951), which evaluates the maximum difference between the empirical cumulative distribution function values of the two data sets.
3.2 Second Stage
The predictive capabilities of climate indexes on the yields of corn (grain), wheat (winter), and soybeans are examined using both traditional statistical approaches and modern machine learning algorithms. In the literature, regression models (Shastry et al. 2017; Shah et al. 2018; Ansarifar et al. 2021) and machine learning techniques (Van Klompenburg et al. 2020; Nigam et al. 2019; Iniyan et al. 2023; Mahesh and Soundrapandiyan 2024) are frequently used for crop yield prediction, primarily due to their simplicity, interpretability, and computational efficiency. In this study, we use GLM and GAM, which are favored for their flexibility regarding distributional assumptions (detailed in Appendix A.1). Furthermore, to demonstrate the advantages of machine learning algorithms over traditional methods, we incorporate two widely used gradient boosting algorithms, XGB and LGBM (detailed in Appendix A.2).
Crop yields are influenced not only by weather conditions during the harvest period but also by extreme weather events throughout the growing season. Such conditions, including temperature extremes and precipitation anomalies, can have significant cumulative effects on crop development and productivity. To capture the full spectrum of potential impacts, all seasonal values for each index are included in the data sets. Additionally, the harvest period of each crop is carefully considered when structuring the data sets. Since corn and soybeans are typically harvested in autumn months, while wheat is harvested in summer months, the seasonal climate indexes are aligned accordingly. For corn and soybeans, the data sets cover the period from winter to autumn, whereas for wheat, they span from autumn to summer for each year. This approach ensures that the data sets encompass the entire year leading up to each crop’s harvest period. The analyses fit a total of 22 models, each incorporating explanatory variables representing various weather measurements and their combinations. A detailed breakdown of each model and the associated data sets is presented in Table 2. Comprehensive information on the data sets, along with the crop yields fitted by each model, is provided in full in Section 4.
Table 2: Explanatory variables included in predicting crop yields
Model
Explanatory
Variables
Model
Explanatory
Variables
Model
Explanatory
Variables
1 CDD 8 T90 15 T90-T10-P-W
2 HDD 9 T10 16 T90-T10-P-D
3 PRE 10 P 17 T90-T10-P-S
4 CDD-HDD 11 T90-T10 18 T90-T10-P-W-D
5 CDD-PRE 12 T90-P 19 T90-T10-P-W-S
6 HDD-PRE 13 T10-P 20 T90-T10-P-D-S
7 CDD-HDD-PRE 14 T90-T10-P 21 T90-T10-P-W-D-S
22 ACI
Note. Based on four meteorological season values for each index.
The choice of explanatory variables for each model follows a logical progression, designed to evaluate the predictive power of various index configurations on crop yields. The first seven models are composed of WBI components and their combinations. The next set of seven models (Models 8–14) features indexes based on ACI components that correspond on a one-to-one basis with WBI variables. Additional ACI components (W, D, S) that are not represented in the initial 14 models are incorporated into the third set of seven models (Models 15–21). Lastly, the final one (Model 22) uses the composite ACI, which aggregates all six components into a single index, offering a comprehensive view of climate conditions in a composite measure.
A flowchart of the proposed methodology is illustrated in Figure 1. The methodology includes two distinct comparison frameworks to evaluate model performances based on different input variables. In the first comparison analysis, the performance of regression models and machine learning algorithms is assessed using principal components derived from the explanatory variables. Since all indexes incorporated into each model are weather-based, they tend to reflect similar meteorological patterns, which may result in multicollinearity issues in regression analyses. To ensure independence between the explanatory variables, and simplify the explanatory data sets without compromising model accuracy, we use PCA and FPCA (see details in Appendix A.3) as dimensionality reduction techniques. By comparing these two techniques, this study also evaluates the performance of FPCA relative to PCA in modeling time-dependent data like climate indexes.
Unlike regression models, machine learning algorithms do not require data simplification or the elimination of multicollinearity among variables. Therefore, in the second comparison analysis, the climate indexes are directly used as input variables in the machine learning algorithms without applying any dimensionality reduction. This approach not only allows for a direct assessment of model performance using the raw climate data but also provides a basis for evaluating the effectiveness of the principal component approach proposed in the methodology.
Figure 1: Flowchart of the proposed methodology
The “scikit-learn” (Fabian 2011) and “scikit-fda” (Ramos-Carreño et al. 2024) Python libraries are used to implement PCA and FPCA with a B-spline basis, respectively. To maintain consistency and simplicity in model execution, the default settings of both packages are applied. From the transformed data sets, the first four principal components ($PC_{1}$, $PC_{2}$, $PC_{3}$, $PC_{4}$), which collectively account for more than 85% of the variance in the data, are selected as the input variables for analyses. This selection ensures that the models capture the majority of the information while reducing dimensionality.
Before conducting the regression analyses, the next critical step involves determining the appropriate link function of GLMs and GAMs, as well as the distribution of crop yield data. The Python package “distfit” (Taskesen 2020) is used to identify the most suitable distribution. This package evaluates 89 univariate distributions and selects the best-fitting one by comparing their residual sum of squares. To ensure consistency across GLMs and GAMs, the log-link function is applied in all models. This function is particularly well-suited for crop yield data, as these values are inherently positive and continuous. The implementation of GLMs and GAMs is carried out using the R packages "stats" (Team 2022) and "mgcv" (Wood 2015), respectively. The GLM used in the analyses is expressed as
$log(E(Y|PC_{1},PC_{2},PC_{3},PC_{4}))=\beta_{0}+\sum_{i=1}^{4}\beta_{i}*PC_{i},$ (9)
whereas the GAM is as follows
$log(E(Y|PC_{1},PC_{2},PC_{3},PC_{4}))=\beta_{0}+\sum_{i=1}^{4}f_{i}(PC_{i}).$ (10)
Here, $Y$ represents the crop yield data, $PC_{i}$ $(i=1,2,3,4)$ denote the principal components obtained through FPCA. $\beta_{i}$’s are the parameters, and $f_{i}(.)$ are smooth functions capturing the potentially nonlinear relationships between principle components and the crop yields. In the analyses, $Y$ and $PC_{i}$ are included as vectors of size $n\times 1$ where $n$ represents the number of observations (years).
[…]
4 Empirical Analysis
4.1 Data
Each component of the ACI consists of a time series, either on a monthly basis, or for meteorological seasons (three months ending February, May, August, November). These time series, starting in 1961, are derived from measurements collected at meteorological stations across the U.S. and Canada. The data are constructed separately for each country, in combination, and across 12 sub-regions. For consistency, this study uses U.S. data, where states are grouped into seven sub-regions based on their geographic locations: Alaska (ALA), Central East Atlantic (CEA), Central West Pacific (CWP), Midwest (MID), Southeast Atlantic (SEA), Southern Plains (SPL), and Southwest Pacific (SWP). The analyses use standardized seasonal values for each of the six ACI components (T90, T10, P, D, W, S) as well as the composite index (ACI). These seasonal data, obtained from the official Actuaries Climate IndexTM website (https://actuariesclimateindex.org/data/, last accessed on April 18, 2025), cover a 63-year period from 1961 to 2023 for all seven sub-regions.
Regarding weather-based indexes, the National Oceanic and Atmospheric Administration (https://www.ncei.noaa.gov/access/monitoring/products/, last accessed on April 18, 2025) provides monthly data on CDD, HDD, and PRE for all U.S. states, covering the period from 1895 to the present. To align with the ACI data, seasonal values for CDD, HDD, and PRE between 1961 and 2023 are calculated by summing the monthly values for each meteorological season. Sub-regional data is subsequently generated by averaging the seasonal values of all states within each region. However, due to the unavailability of CDD and HDD data for Alaska, the analyses are limited to the six remaining regions of the U.S.
Annual crop yield data (in bushels per acre) for three major agricultural products-corn (grain), wheat (winter), and soybeans-produced in the U.S. is obtained from the United States Department of Agriculture (https://quickstats.nass.usda.gov/, last accessed on April 18, 2025) for all states, covering the period from 1961 to 2023. Following the same approach used in regional weather-based index calculations, sub-regional crop yields are calculated by averaging the annual yields of the states within each region. Analyses for corn and wheat are conducted across all regions, while for soybeans, due to data limitations, the analyses are restricted to four regions: CEA, CWP, MID, and SEA.
4.2 Results
The analyses start with determining the dependence among three groups (Table 1). It is noticed that all correlation coefficients, $r$, are statistically significant ($p$-value $<0.01$), as shown in Table 3. Despite methodological differences in constructing the indexes, strong positive correlations are observed between the weather-based indexes and their corresponding components of the ACI. To assess the similarity of the distributions, we run K-S test which evaluates the null hypothesis that the distributions are identical. Table 3 also presents the indexes that exhibit statistically similar distributions, as indicated by $p$-values greater than 0.05. The results reveal that the precipitation-related indexes in Group-3 exhibit similar distributions across most regions compared to temperature measurements analyzed in the first two groups. Detailed descriptive statistics for each group are provided in Table B1 of the Appendix. Additionally, the indexes and their corresponding cumulative distribution function (CDF) values are illustrated in Figures B1 and B2 of the Appendix.
Table 3: Similarity tests within each group
Region
CEA CWP MID SEA SPL SWP
Group-1 (CDD-T90) r 0.88 0.90 0.89 0.94 0.86 0.96
K-S test (p) 0.089 <0.01 <0.01 <0.01 <0.01 0.408
Group-2 (HDD-T10) r 0.89 0.83 0.86 0.87 0.81 0.79
K-S test (p) 0.011 <0.01 <0.01 0.019 0.204 0.055
Group-3 (PRE-P) r 0.83 0.81 0.87 0.87 0.68 0.85
K-S test (p) 0.293 0.293 0.408 0.089 <0.01 0.089
Note. r: Spearman rank correlation coefficient.
In the second stage of the analysis, explanatory variables in each model are transformed into their principal components using PCA and FPCA. Figure 2 displays the total variation in the data sets, represented as regional averages, explained by the first four principal components derived from both methods for each model. The results demonstrate that the total variance explained by FPCA consistently surpasses that of PCA across all models, highlighting FPCA’s superior performance when dealing with time-dependent data sets like climate indexes.
Figure 2: Explained variance comparisons of PCA and FPCA as regional averages
Upon closer examination of Figure 2, the models (1, 2, 3, 8, 9, 10, and 22) show a total explained variance of 100% for both PCA and FPCA. This is expected, as these models contain only one explanatory variable represented by its four seasonal values. Consequently, PCA and FPCA serve purely as data transformation tools for these cases without affecting the total variance explained.
For a detailed comparison of ACI components and their WBI counterparts, it is important to recall the structure of models outlined in Table 2. The first seven models are based on WBI components, while the subsequent seven models (Models 8–14) use their ACI counterparts. Comparing the total explained variance for paired models (e.g., 1 vs. 8, 2 vs. 9, 3 vs. 10, and so forth), it is clear that these values are nearly equivalent. For instance, the total variance explained by PCA for Model 4 is 83%, whereas it is 84% for Model 11. Similarly, FPCA explains 88% of the variance for Model 7 and 90% for Model 14. Since the choice of explanatory variables directly influence regression outcomes, ensuring high and comparable explained variance values across both index groups is crucial. Given that FPCA consistently explains a higher proportion of variance across all models, with values exceeding 85%, the principal components derived from FPCA are selected for the subsequent analyses, ensuring better model accuracy and reliability.
The distributions from the exponential family that best fit each crop yield for each region are identified, and the results are presented in Table 4. These distributions form the basis for estimating crop yields using GLMs and GAMs. Moreover, the incorporation of region-specific distributions ensures that the regression models account for spatial variability in crop yields, further improving the precision and relevance of the analyses.
Table 4: Best-fitting distributions for each crop
Region Crop
Corn (Grain) Wheat (Winter) Soybeans
CEA gamma gamma gamma
CWP normal normal -
MID normal gamma normal
SEA gamma gamma gamma
SPL normal normal gamma
SWP normal normal -
[…]
Method Model
1 2 3 4 5 6 7 8 9 10 11
GLM 0.288 0.296 0.170 0.286 0.334 0.353 0.384 0.217 0.231 0.292 0.248
GAM 0.384 0.388 0.309 0.396 0.475 0.414 0.444 0.389 0.446 0.416 0.381
XGB 0.738 0.747 0.691 0.753 0.757 0.766 0.769 0.690 0.732 0.734 0.708
LGBM 0.327 0.309 0.219 0.352 0.349 0.406 0.416 0.263 0.329 0.339 0.315
12 13 14 15 16 17 18 19 20 21 22
GLM 0.372 0.391 0.406 0.392 0.378 0.574 0.420 0.584 0.551 0.515 0.403
GAM 0.481 0.553 0.454 0.457 0.533 0.644 0.469 0.689 0.677 0.625 0.516
XGB 0.775 0.797 0.806 0.780 0.790 0.858 0.790 0.866 0.864 0.856 0.788
LGBM 0.437 0.456 0.427 0.401 0.429 0.571 0.414 0.616 0.581 0.523 0.406
Figure 3: MAPE values (Test) of each model as regional averages, Corn (Grain)
[…]
Method Model
1 2 3 4 5 6 7 8 9 10 11
GLM 0.262 0.296 0.118 0.301 0.282 0.345 0.333 0.215 0.260 0.239 0.242
GAM 0.439 0.515 0.360 0.527 0.372 0.400 0.433 0.398 0.436 0.397 0.359
XGB 0.725 0.756 0.659 0.752 0.730 0.737 0.776 0.725 0.738 0.696 0.723
LGBM 0.289 0.353 0.162 0.346 0.300 0.321 0.320 0.328 0.318 0.260 0.288
12 13 14 15 16 17 18 19 20 21 22
GLM 0.360 0.349 0.355 0.397 0.333 0.526 0.366 0.552 0.532 0.494 0.439
GAM 0.493 0.529 0.458 0.516 0.483 0.599 0.439 0.619 0.593 0.538 0.535
XGB 0.781 0.741 0.747 0.786 0.765 0.842 0.786 0.842 0.841 0.825 0.788
LGBM 0.367 0.371 0.327 0.417 0.362 0.524 0.408 0.552 0.538 0.507 0.440
Figure 4: MAPE values (Test) of each model as regional averages, Wheat (Winter)
[…]
Method Model
1 2 3 4 5 6 7 8 9 10 11
GLM 0.250 0.265 0.254 0.234 0.377 0.387 0.442 0.128 0.182 0.319 0.164
GAM 0.327 0.350 0.421 0.352 0.491 0.426 0.534 0.302 0.515 0.431 0.318
XGB 0.675 0.723 0.706 0.689 0.774 0.773 0.784 0.666 0.727 0.732 0.663
LGBM 0.222 0.252 0.258 0.286 0.357 0.426 0.425 0.146 0.314 0.326 0.235
12 13 14 15 16 17 18 19 20 21 22
GLM 0.341 0.378 0.402 0.362 0.389 0.602 0.374 0.597 0.573 0.567 0.381
GAM 0.467 0.571 0.479 0.441 0.615 0.721 0.422 0.641 0.695 0.606 0.528
XGB 0.740 0.786 0.761 0.770 0.768 0.862 0.752 0.861 0.847 0.838 0.756
LGBM 0.367 0.436 0.354 0.368 0.387 0.558 0.348 0.537 0.510 0.526 0.343
Figure 5: MAPE values (Test) of each model as regional averages, Soybeans
When analyzing the best-performing models for the three crops (Model 19 for corn and wheat, Model 17 for soybeans), it becomes clear that, in addition to temperature and precipitation variables, sea-level change plays a significant role in predicting yields. One possible explanation for this is that variations in sea-level can impact crop production more broadly, particularly through mechanisms such as increased soil salinity, flooding, and the loss of agricultural land. Furthermore, wind appears to have a less pronounced effect on soybean yield compared to corn and wheat. This could be attributed to the relatively short stature of soybean plants, making them more resistant to wind-induced damage. In contrast, taller crops like corn and wheat may be more vulnerable to lodging [^1] Lodging is the bending over of the stems of crops which makes them very difficult to harvest. It usually happens due to the wind-related stress. and other wind-related stresses, which can negatively impact their overall yield.
[…]
Another critical observation is that the final model (Model 22), which uses the composite ACI by aggregating all six components into a single index, consistently underperforms compared to Model 21, which incorporates each component separately, across all crops and methods. These findings highlight that the composite ACI index alone fails to fully capture and explain the complex weather conditions affecting crop yields. This suggests that treating the individual components separately provides a more nuanced and effective modeling approach, ensuring greater accuracy in predicting yield variations.
The second comparison analysis is designed to evaluate the performance sensitivity of machine learning algorithms with respect to different input variables. Table 8 presents only the results of XGB for the second comparison analysis, as it outperformed LGBM in earlier analyses. In Table 8, XGB refers to the results obtained using principal components derived from climate indexes, while XGB-C corresponds to the scenario where the climate indexes are directly used as input variables. Notably, with the XGB-C method, Model 21 (which incorporates all six components of ACI and thus constitutes the largest data set) consistently yields the best performance, with only a slight difference compared to the second-best models. Among the remaining models that exclude one or more components, the second-best performing models remain consistent with those from the first comparison analysis (Model 19 for corn and wheat, Model 17 for soybeans). One possible explanation for this is that the direct use of climate indexes preserves all original information in the data sets, enabling machine learning algorithms to process and use a greater number of features more effectively, compared to the more compressed data sets derived through principal component analysis. Regional performance metrics of XGB-C are given in Tables B11B11 of the Appendix.
It is also noteworthy that across all analyses, the best-performing models for each crop remain consistent across all methods (GLM, GAM, XGB, LGBM). This consistency underscores the substantial impact of specific climate indexes on crop yields, demonstrating that their predictive value holds true regardless of the modeling approach.
Table 8: Performance Comparison-2 as regional averages for each crop
Corn (Grain) Wheat (Winter) Soybeans
pR2 (Train) MAPE (Test) pR2 (Train) MAPE (Test) pR2 (Train) MAPE (Test)
Model XGB XGB-C XGB XGB-C XGB XGB-C XGB XGB-C XGB XGB-C XGB XGB-C
1 0.738 0.741 0.218 0.199 0.725 0.704 0.170 0.166 0.675 0.713 0.213 0.226
2 0.747 0.723 0.240 0.243 0.756 0.746 0.226 0.197 0.723 0.703 0.260 0.276
3 0.691 0.663 0.321 0.323 0.659 0.660 0.269 0.267 0.706 0.723 0.289 0.285
4 0.753 0.822 0.226 0.221 0.752 0.814 0.200 0.187 0.689 0.819 0.247 0.264
5 0.757 0.848 0.204 0.207 0.730 0.822 0.200 0.190 0.774 0.834 0.201 0.222
6 0.766 0.842 0.263 0.246 0.737 0.849 0.217 0.193 0.773 0.859 0.268 0.260
7 0.769 0.877 0.212 0.208 0.776 0.881 0.184 0.184 0.784 0.886 0.205 0.233
8 0.690 0.726 0.242 0.260 0.725 0.718 0.220 0.209 0.666 0.630 0.273 0.294
9 0.732 0.728 0.239 0.235 0.738 0.755 0.214 0.202 0.727 0.731 0.258 0.253
10 0.734 0.750 0.310 0.326 0.696 0.740 0.227 0.243 0.732 0.750 0.287 0.303
11 0.708 0.811 0.256 0.218 0.723 0.811 0.246 0.208 0.663 0.776 0.265 0.236
12 0.775 0.880 0.224 0.235 0.781 0.852 0.208 0.201 0.740 0.845 0.250 0.260
13 0.797 0.853 0.217 0.225 0.741 0.859 0.209 0.186 0.786 0.846 0.206 0.245
14 0.806 0.894 0.243 0.206 0.747 0.885 0.237 0.187 0.761 0.879 0.274 0.234
15 0.780 0.933 0.196 0.179 0.786 0.922 0.161 0.142 0.770 0.927 0.214 0.209
16 0.790 0.909 0.261 0.211 0.765 0.899 0.232 0.193 0.768 0.893 0.249 0.230
17 0.858 0.950 0.145 0.141 0.842 0.943 0.152 0.132 0.862 0.957ϕ 0.134 0.134ϕ
18 0.790 0.941 0.201 0.180 0.786 0.935 0.192 0.138 0.752 0.935 0.228 0.197
19 0.866 0.963ϕ 0.143 0.136ϕ 0.842 0.954ϕ 0.146 0.124ϕ 0.861 0.941 0.142 0.134
20 0.864 0.961 0.161 0.144 0.841 0.949 0.184 0.133 0.847 0.953 0.139 0.140
21 0.856 0.964 0.150 0.135 0.825 0.960 0.161 0.119 0.838 0.963 0.148 0.130
22 0.788 0.780 0.203 0.203 0.788 0.773 0.197 0.166 0.756 0.772 0.249 0.228
Note. Best results, ϕSecond best results.
5 Concluding Remarks
This study examines the Actuaries Climate IndexTM within the agricultural sector and compares its effectiveness with widely recognized weather-based indexes. The focus is on evaluating the predictive capabilities of both ACI components and traditional weather-based indexes in predicting three selected crop yields. By fitting 22 distinct models, each constructed with different climate indexes and their combinations, this paper also investigates the individual contribution of each weather index to model performance. Based on high goodness-of-fit test results and low MAPE values, climate indexes exhibit strong predictive power in modelling crop yields. The key insights drawn from the analyses can be summarized as follows:
  1. (i)
    The FPCA consistently outperforms the traditional PCA in terms of explanatory variance, when applied to time-dependent climate indexes.
  2. (ii)
    The four principal components derived through FPCA are sufficiently effective in capturing the underlying patterns in crop yield variations.
  3. (iii)
    ACI components and their WBI counterparts exhibit strong correlation, and deliver similar predictive performance in predicting crop yields, especially with machine learning algorithms.
  4. (iv)
    The inclusion of indexes representing additional weather conditions—such as wind speed, and sea-level changes—alongside temperature and precipitation, significantly improves model performance.
  5. (v)
    The composite ACI index alone fails to fully capture and explain the complex weather conditions affecting crop yields.
The outcomes suggest that the ACI framework has a potential to serve as a valuable climate risk indicator not only for the agricultural sector but also for the finance and insurance industries, where climate-related metrics are becoming increasingly critical for risk assessment and decision-making.
Despite the promising performance metrics achieved, the study can be further improved in several aspects. One notable limitation is that regional crop yields were assessed by averaging the annual yields of the states within each region. This approach overlooks the fact that some states may have a greater influence on regional agricultural output due to their size or productivity levels. Moreover, the selection of the best-performing models for each crop is based on regional average results, which may not accurately reflect optimal performance at the regional level. Consequently, identifying and evaluating the best models for each specific region or state would enhance the precision and applicability of the findings, particularly for targeted applications in the insurance sector.
The proposed methodology also contributes to the literature on forecasting crop yields. Forecasting weather-related factors that significantly impact crop yields such as temperature, precipitation, and sea-level, often requires complex, time-consuming methods due to their inherently dynamic nature. However, by incorporating future projections of components derived through the FPCA method, crop yield forecasts can be achieved in a more efficient and convenient manner.
Acknowledgments
Jose Garrido gratefully acknowledges the financial support of the Natural Sciences and Engineering Research Council of Canada (RGPIN-2017-06643).
Appendix
Appendix A Preliminaries
In this section, we provide an overview of generalized linear models (GLMs) and generalized additive models (GAMs), which belong to the class of generalized statistical models, as well as Extreme Gradient Boosting (XGB) and Light Gradient Boosting Machine (LGBM), which are classified as machine learning algorithms. Furthermore, we review principal component analysis (PCA) and functional principal component analysis (FPCA), two dimensionality reduction techniques used to preprocess and simplify data sets before being incorporated into the models.
A.1 Generalized Statistical Models
GLMs are widely used regression framework that extends traditional linear models by accommodating response variables with error distributions other than normal (Nelder and Wedderburn 1972). It is applied in various fields due to its flexibility. Simply, the model setting consists of three key components: a distribution, a linear predictor, and a link function. The general formulation of GLMs is:
$g(E(Y|X))=X\beta,$ (11)
where $Y$ represents the response variable, chosen from the exponential family (e.g., Binomial, Poisson, or Gaussian). The linear predictor can be expressed as $\eta=X\beta$, where $X$ is the matrix of independent variables, and $\beta$ is the coefficient vector. The link function $g(.)$ connects the conditional expectation of $Y$ to the linear predictor. Commonly used link functions, depending on the chosen distribution, are the identity, logarithmic and logistic functions.
While the GLMs are powerful and widely used tool for estimation and prediction, they have some limitations. One significant drawback is their assumption of a linear relationship between the predictors and the response variable. When the true relationship is non-linear, the results of GLMs can become biased or yield inaccurate predictions. GAMs address this limitation by extending the GLMs framework to accommodate non-linear relationships (Hastie 2017; Wood 2017). GAMs relax the linearity assumption by replacing linear terms with smooth functions, enabling each predictor to have its own potentially non-linear effect on the response variable. In general, the GAMs take the form:
$g(E(Y|X))=\eta=\beta_{0}+f_{1}(X_{1})+f_{2}(X_{2})+\dots+f_{p}(X_{p}),$ (12)
where $\beta_{0}$ is the intercept term, and $f_{j}(X_{j})$ $(j=1,\dots,p)$ are smooth functions that model the non-linear effects of the predictors $X_{j}$ on the response variable. The flexibility of GAMs allow them to capture complex relationships while retaining interpretability, making it an effective tool for modeling real-world problems with non-linear dependencies.
A.2 Machine Learning Algorithms
XGB is an advanced implementation of the gradient boosting framework, optimized for performance and computational efficiency (Chen and Guestrin 2016). It is a well-optimized and scalable algorithm that is particularly suitable for regression, classification (binary and multiclass), and ranking problems. The algorithm functions by adding weak learners, such as decision trees, to the ensemble iteratively, with each successive model trying to correct the errors of the previous models. It uses a regularized objective function that includes both the training loss and a penalty term to control model complexity and avoid overfitting. The regularized objective can be written as:
$\mathcal{L}^{(t)}=\sum_{i=1}^{n}\ell\left(y_{i},\hat{y}_{i}^{(t-1)}+f_{t}(x_{i})\right)+\Omega(f_{t}),\quad\text{where}\quad\Omega(f)=\gamma T+\frac{1}{2}\lambda\|\mathbf{w}\|^{2}.$ (13)
Here, $\ell$ denotes a differentiable loss function, $\hat{y}_{i}^{(t-1)}$ is the prediction at iteration $t-1$, $\Omega(f_{t})$ penalizes the complexity of each tree $f_{t}$, with $T$ being the number of leaves and $\mathbf{w}$ the vector of leaf weights.
On the other hand, LGBM is a gradient boosting framework designed for speed and efficiency, especially on large-scale data sets with high dimensionality (Ke et al. 2017). It uses Gradient-based One-Side Sampling which selectively retains instances with large gradients to better capture under-fitted samples. Unlike level-wise growth used in other tree-based models, LGBM adopts a leaf-wise tree growth strategy that expands the leaf with the largest loss reduction, which often leads to lower loss with fewer iterations. The model minimizes the following objective function:
$L(\theta)=\sum_{i=1}^{n}\ell(y_{i},\hat{y}_{i})+\sum_{j=1}^{K}\Omega(f_{j}),$ (14)
where $L(\theta)$ is the total loss, $n$ is the number of samples, $K$ is the number of trees, and $\Omega(f_{j})$ is a regularization term that penalizes the complexity of the model to avoid overfitting.
A.3 Dimensionality Reduction in Modeling
For high dimensional explanatory variables, it is important to be able to reduce the number of variables while preserving valuable information as best possible. Simplifying the data set enhances interpretability, reduces computational complexity, and minimizes redundancy and noise. One widely used method for dimension reduction is principal component analysis. PCA transforms the original variables into a new set of uncorrelated variables called principal components (Pearson 1901; Hotelling 1933). Standardization is needed when the variables in the data set are on different scales.
Given a data matrix $X$ of size $n\times p$ where $n$ represents the number of observations and $p$ is the number of variables, each element in $X$ can be standardized around its mean value, normalized by its standard deviation, thus obtaining a matrix of $Z$. PCA seeks to find the principal components by calculating the eigenvectors and eigenvalues of the covariance matrix, $\Sigma_{p\times p}$, given by,
$\Sigma=\frac{1}{n-1}Z^{T}Z.$ (15)
Singular value decomposition (SDV) is then applied to decompose the covariance matrix, $\Sigma$. The eigenvalue equation is expressed as:
$\Sigma V=\lambda V,$ (16)
where $\lambda_{p\times p}$ is a diagonal matrix consisting of the set of all eigenvalues of $\Sigma$ along its principal diagonal, and zero for all other elements, and $V_{p\times p}$ is the matrix consisting of all eigenvectors of $\Sigma$, one eigenvector per column. The component with the largest eigenvalue becomes the first principal component which explains the greatest variation in the data. The explained variance, $\sigma^{2}_{EXP}$, by first $k$ eigenvalues is then calculated by:
$\sigma^{2}_{EXP}=\frac{\sum_{j=1}^{k}\lambda_{j}}{\sum_{j=1}^{p}\lambda_{j}}.$ (17)
The order of the data points in the data set analyzed by PCA has no impact on the results. This is because PCA treats the data set as a collection of vectors and does not account for the sequential or spatial arrangement of the data; the order can be permuted without altering the outcome. This makes PCA unsuitable for analyzing functional data, such as time series or curves, where the sequence of observations holds critical information. Functional principal component analysis extends traditional PCA, tailoring it to analyze data where each observation is a continuous function defined over some domain, rather than a discrete vector (Ramsay and Silverman 2002; Shang 2014).
Let $F(t)$ denote functional data defined over a domain $t\in[a,b]$, where $t$ represents time. Then, functional data can be represented using a basis function expansion, expressed as:
$F(t)=\sum_{k=1}^{\infty}\beta_{k}\phi_{k}(t)+\epsilon(t),$ (18)
where $\phi_{k}(t)$ are the eigenfunctions (e.g., B-splines, Fourier or monomial basis), $\beta_{k}$ are the coefficients associated with each functional principal component, and $\epsilon(t)$ represents noise. Instead of a covariance matrix, FPCA uses a covariance function $C(s,t)$ for decomposition. The eigenfunctions $\phi_{k}(t)$, and their corresponding eigenvalues $\lambda_{k}$ are found by solving the functional eigenvalue problem:
[…]
Appendix B Tables and Figures
Table B1: Descriptive statistics of each group
n = 63 Region
CEA CWP MID SEA SPL SWP
Group-1 CDD Mean ± SD 504 ± 96 243 ± 72 617 ± 109 1175 ± 109 794 ± 90 888 ± 100
Min. 326 84 356 928 526 697
Max. 703 427 839 1464 1016 1086
T90 Mean ± SD 0.40 ± 1.20 0.49 ± 1.31 0.01 ± 0.89 0.54 ± 1.21 0.54 ± 1.42 1.06 ± 1.69
Min. -1.48 -1.65 -1.32 -1.25 -1.61 -1.91
Max. 3.57 5.15 3.57 4.49 5.32 5.13
Group-2 HDD Mean ± SD 3274 ± 275 2727 ± 184 3658 ± 326 1877 ± 236 3346 ± 269 2375 ± 166
Min. 2715 2285 3002 1415 2797 2024
Max. 3832 3237 4382 2452 4208 2799
T10 Mean ± SD -0.46 ± 1.05 -0.15 ± 0.89 -0.30 ± 0.96 -0.44 ± 0.98 -0.28 ± 0.93 -0.31 ± 0.88
Min. -2.11 -1.50 -1.81 -1.84 -1.86 -1.84
Max. 1.98 2.42 2.41 2.49 2.88 2.51
Group-3 PRE Mean ± SD 11.29 ± 2.21 8.31 ± 1.86 10.07 ± 1.73 13.48 ± 2.61 7.02 ± 1.19 3.61 ± 1.04
Min. 6.79 3.88 6.10 7.49 4.49 1.54
Max. 19.04 13.49 13.82 10.21 10.21 7.02
P Mean ± SD 0.10 ± 1.00 0.58 ± 1.16 0.26 ± 1.03 -0.02 ± 0.94 0.33 ± 1.39 0.19 ± 0.97
Min. -2.05 -1.43 -2.13 -2.07 -1.65 -1.92
Max. 3.48 2.85 2.95 1.85 5.27 2.93
Note. n: Number of observations (years), SD: Standard deviation, Min: Minimum, Max: Maximum.
(a) Group-1
(b) Group-2
(c) Group-3
Figure B1: Group indexes of each region over time
(a) Group-1
(b) Group-2
(c) Group-3
Figure B2: CDF plots of group indexes for each region
[…]
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.174 0.276 0.345 0.197 0.493 0.466 0.458 0.122 0.251 0.340 0.347 0.397 0.417 0.374 0.421 0.344 0.632 0.431 0.608 0.608 0.604 0.428
CWP 0.297 0.292 0.182 0.319 0.240 0.391 0.361 0.313 0.282 0.502 0.322 0.552 0.565 0.593 0.538 0.493 0.452 0.666 0.549 0.504 0.303 0.328
MID 0.228 0.236 0.249 0.184 0.336 0.291 0.312 0.112 0.101 0.312 0.022 0.322 0.297 0.307 0.338 0.237 0.307 0.292 0.338 0.237 0.292 0.156
SEA 0.295 0.334 0.113 0.292 0.260 0.280 0.342 0.172 0.214 0.198 0.195 0.185 0.312 0.311 0.231 0.292 0.796 0.220 0.751 0.704 0.687 0.520
SPL 0.251 0.198 0.105 0.224 0.193 0.212 0.299 0.130 0.128 0.280 0.120 0.286 0.239 0.306 0.337 0.381 0.755 0.386 0.721 0.734 0.736 0.533
SWP 0.481 0.441 0.026 0.499 0.481 0.479 0.533 0.455 0.412 0.121 0.486 0.489 0.516 0.542 0.487 0.523 0.504 0.524 0.538 0.519 0.466 0.452
Wheat (Winter) CEA 0.328 0.473 0.065 0.486 0.308 0.440 0.414 0.274 0.444 0.122 0.338 0.334 0.368 0.378 0.484 0.406 0.689 0.478 0.621 0.668 0.674 0.655
CWP 0.125 0.257 0.216 0.160 0.226 0.390 0.248 0.231 0.270 0.567 0.281 0.661 0.646 0.626 0.581 0.483 0.491 0.487 0.517 0.486 0.290 0.321
MID 0.202 0.234 0.117 0.284 0.220 0.266 0.281 0.102 0.123 0.159 0.128 0.152 0.190 0.198 0.359 0.262 0.198 0.305 0.359 0.262 0.305 0.174
SEA 0.295 0.283 0.067 0.271 0.193 0.223 0.242 0.227 0.223 0.111 0.235 0.178 0.111 0.122 0.241 0.117 0.643 0.195 0.606 0.631 0.578 0.570
SPL 0.281 0.200 0.203 0.277 0.361 0.358 0.387 0.141 0.201 0.308 0.128 0.341 0.377 0.365 0.280 0.324 0.735 0.306 0.774 0.741 0.727 0.528
SWP 0.341 0.326 0.038 0.330 0.382 0.394 0.423 0.317 0.300 0.167 0.340 0.492 0.404 0.442 0.435 0.403 0.398 0.426 0.433 0.403 0.387 0.384
Soybeans CEA 0.201 0.328 0.368 0.256 0.542 0.565 0.553 0.167 0.285 0.332 0.375 0.434 0.440 0.430 0.387 0.438 0.614 0.437 0.597 0.624 0.596 0.462
MID 0.281 0.282 0.255 0.250 0.391 0.362 0.416 0.094 0.152 0.336 0.052 0.346 0.345 0.366 0.389 0.326 0.366 0.344 0.389 0.326 0.344 0.148
SEA 0.231 0.240 0.248 0.196 0.318 0.395 0.419 0.096 0.149 0.284 0.130 0.193 0.384 0.392 0.285 0.316 0.664 0.259 0.666 0.600 0.572 0.390
SPL 0.285 0.210 0.146 0.234 0.258 0.227 0.382 0.154 0.145 0.325 0.099 0.390 0.342 0.422 0.387 0.477 0.763 0.457 0.739 0.742 0.753 0.525
Table B3: MAPE values of GLMs for each crop and region (Test Set)
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.241 0.224 0.288 0.234 0.214 0.204 0.201 0.318 0.225 0.317 0.258 0.309 0.244 0.274 0.211 0.232 0.085 0.176 0.104 0.114 0.134 0.122
CWP 0.129 0.251 0.351 0.254 0.229 0.314 0.243 0.209 0.277 0.288 0.238 0.244 0.278 0.240 0.243 0.284 0.301 0.164 0.147 0.268 0.226 0.319
MID 0.281 0.329 0.301 0.321 0.251 0.314 0.255 0.340 0.350 0.305 0.367 0.292 0.304 0.293 0.261 0.275 0.293 0.279 0.261 0.275 0.347 0.279
SEA 0.168 0.263 0.399 0.269 0.243 0.330 0.238 0.310 0.302 0.339 0.337 0.262 0.227 0.235 0.222 0.246 0.323 0.234 0.347 0.312 0.146 0.321
SPL 0.159 0.232 0.250 0.244 0.187 0.264 0.184 0.151 0.241 0.248 0.190 0.168 0.235 0.208 0.233 0.181 0.261 0.229 0.168 0.246 0.068 0.291
SWP 0.176 0.110 0.214 0.124 0.108 0.120 0.051 0.157 0.107 0.225 0.136 0.124 0.102 0.057 0.069 0.102 0.058 0.072 0.049 0.063 0.159 0.077
Wheat (Winter) CEA 0.174 0.182 0.316 0.162 0.201 0.198 0.176 0.248 0.186 0.307 0.235 0.216 0.235 0.231 0.178 0.211 0.033 0.162 0.045 0.059 0.041 0.146
CWP 0.218 0.213 0.220 0.181 0.150 0.225 0.181 0.198 0.226 0.140 0.220 0.138 0.154 0.119 0.112 0.174 0.171 0.187 0.170 0.144 0.201 0.161
MID 0.254 0.290 0.283 0.266 0.242 0.269 0.248 0.317 0.310 0.285 0.325 0.293 0.302 0.291 0.282 0.273 0.291 0.274 0.273 0.282 0.274 0.307
SEA 0.159 0.220 0.346 0.158 0.233 0.269 0.233 0.247 0.225 0.295 0.269 0.274 0.286 0.276 0.165 0.267 0.106 0.216 0.072 0.081 0.085 0.107
SPL 0.177 0.212 0.210 0.183 0.155 0.205 0.175 0.146 0.213 0.214 0.172 0.184 0.213 0.195 0.208 0.197 0.127 0.210 0.111 0.110 0.154 0.093
SWP 0.144 0.114 0.061 0.133 0.073 0.067 0.063 0.167 0.086 0.071 0.124 0.121 0.037 0.073 0.051 0.055 0.107 0.043 0.055 0.068 0.078 0.220
Soybeans CEA 0.193 0.178 0.255 0.185 0.163 0.153 0.154 0.271 0.179 0.287 0.207 0.264 0.203 0.225 0.187 0.170 0.100 0.143 0.100 0.111 0.107 0.093
MID 0.232 0.283 0.280 0.273 0.215 0.275 0.223 0.307 0.304 0.279 0.318 0.268 0.266 0.270 0.240 0.239 0.240 0.243 0.270 0.239 0.243 0.308
SEA 0.231 0.289 0.319 0.294 0.231 0.276 0.238 0.319 0.309 0.285 0.324 0.263 0.218 0.236 0.221 0.244 0.093 0.229 0.060 0.088 0.106 0.216
SPL 0.184 0.226 0.229 0.221 0.157 0.250 0.175 0.175 0.232 0.231 0.183 0.166 0.212 0.184 0.195 0.153 0.108 0.196 0.114 0.103 0.146 0.129
[…]
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.199 0.415 0.345 0.336 0.493 0.528 0.498 0.195 0.650 0.547 0.461 0.596 0.565 0.374 0.623 0.446 0.670 0.473 0.704 0.618 0.712 0.673
CWP 0.548 0.391 0.452 0.470 0.560 0.391 0.362 0.455 0.480 0.616 0.442 0.673 0.638 0.635 0.619 0.558 0.534 0.681 0.573 0.620 0.435 0.375
MID 0.228 0.236 0.427 0.215 0.336 0.332 0.490 0.193 0.338 0.442 0.224 0.517 0.616 0.365 0.405 0.482 0.365 0.292 0.482 0.405 0.292 0.156
SEA 0.328 0.497 0.387 0.516 0.476 0.334 0.412 0.563 0.443 0.334 0.207 0.185 0.493 0.366 0.269 0.591 0.796 0.312 0.729 0.759 0.778 0.693
SPL 0.393 0.201 0.214 0.281 0.313 0.271 0.299 0.413 0.310 0.345 0.205 0.294 0.426 0.330 0.343 0.465 0.877 0.426 0.893 0.919 0.844 0.703
SWP 0.612 0.589 0.026 0.558 0.675 0.630 0.602 0.514 0.453 0.212 0.750 0.621 0.580 0.653 0.487 0.653 0.620 0.630 0.756 0.741 0.689 0.497
Wheat (Winter) CEA 0.468 0.589 0.418 0.608 0.308 0.502 0.496 0.349 0.539 0.465 0.338 0.429 0.634 0.455 0.681 0.583 0.639 0.559 0.705 0.724 0.767 0.677
CWP 0.475 0.753 0.216 0.621 0.327 0.390 0.478 0.595 0.492 0.635 0.453 0.811 0.700 0.698 0.646 0.559 0.541 0.496 0.498 0.533 0.392 0.349
MID 0.372 0.411 0.622 0.368 0.351 0.357 0.281 0.279 0.532 0.281 0.252 0.311 0.296 0.364 0.461 0.432 0.461 0.328 0.364 0.432 0.437 0.328
SEA 0.506 0.423 0.354 0.667 0.287 0.274 0.415 0.451 0.361 0.184 0.360 0.527 0.354 0.272 0.328 0.189 0.667 0.393 0.832 0.652 0.669 0.588
SPL 0.386 0.453 0.267 0.355 0.579 0.390 0.463 0.396 0.366 0.428 0.315 0.341 0.608 0.375 0.280 0.423 0.800 0.325 0.792 0.757 0.570 0.734
SWP 0.429 0.459 0.284 0.540 0.382 0.485 0.463 0.317 0.325 0.389 0.438 0.536 0.584 0.585 0.698 0.712 0.487 0.530 0.525 0.463 0.390 0.532
Soybeans CEA 0.201 0.427 0.387 0.423 0.595 0.644 0.579 0.167 0.608 0.559 0.447 0.643 0.597 0.430 0.545 0.534 0.718 0.453 0.602 0.682 0.652 0.573
MID 0.323 0.375 0.494 0.250 0.397 0.396 0.548 0.406 0.672 0.360 0.277 0.593 0.620 0.584 0.435 0.616 0.616 0.344 0.435 0.584 0.344 0.423
SEA 0.358 0.375 0.469 0.465 0.562 0.395 0.488 0.383 0.364 0.443 0.135 0.193 0.559 0.479 0.383 0.728 0.723 0.309 0.685 0.729 0.659 0.472
SPL 0.426 0.223 0.334 0.272 0.408 0.268 0.520 0.249 0.414 0.362 0.413 0.440 0.507 0.423 0.400 0.580 0.826 0.582 0.842 0.786 0.768 0.642
Table B5: MAPE values of GAMs for each crop and region (Test Set)
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.228 0.259 0.288 0.289 0.214 0.210 0.185 0.302 0.263 0.303 0.259 0.313 0.255 0.274 0.201 0.289 0.075 0.163 0.102 0.131 0.224 0.108
CWP 0.294 0.289 0.313 0.400 0.562 0.314 0.243 0.402 0.230 0.305 0.290 0.256 0.261 0.246 0.215 0.288 0.241 0.165 0.172 0.243 0.190 0.302
MID 0.281 0.329 0.332 0.314 0.251 0.304 0.295 0.359 0.339 0.332 0.389 0.310 0.307 0.305 0.234 0.228 0.305 0.279 0.234 0.228 0.347 0.279
SEA 0.138 0.248 0.459 0.214 0.254 0.335 0.226 0.295 0.272 0.364 0.330 0.262 0.214 0.228 0.201 0.450 0.323 0.280 0.272 0.183 0.194 0.175
SPL 0.129 0.230 0.275 0.238 0.156 0.264 0.184 0.170 0.222 0.259 0.171 0.164 0.178 0.196 0.231 0.172 0.085 0.226 0.079 0.167 0.118 0.151
SWP 0.138 0.104 0.214 0.105 0.159 0.178 0.068 0.097 0.103 0.221 0.260 0.228 0.139 0.155 0.069 0.118 0.072 0.130 0.114 0.228 0.129 0.368
Wheat (Winter) CEA 0.172 0.197 0.354 0.176 0.200 0.205 0.161 0.235 0.195 0.239 0.235 0.210 0.240 0.239 0.150 0.269 0.042 0.165 0.081 0.088 0.040 0.148
CWP 0.456 0.207 0.220 0.187 0.168 0.225 0.175 0.165 0.257 0.148 0.240 0.443 0.160 0.117 0.090 0.163 0.168 0.186 0.134 0.170 0.198 0.160
MID 0.188 0.294 0.293 0.288 0.226 0.265 0.248 0.382 0.286 0.265 0.341 0.320 0.302 0.311 0.281 0.340 0.311 0.282 0.281 0.340 0.282 0.363
SEA 0.229 0.222 0.399 0.380 0.271 0.275 0.262 0.319 0.204 0.290 0.228 0.305 0.287 0.229 0.150 0.307 0.104 0.225 0.074 0.085 0.066 0.132
SPL 0.171 0.238 0.217 0.212 0.139 0.199 0.174 0.145 0.175 0.228 0.168 0.184 0.181 0.192 0.208 0.201 0.131 0.215 0.109 0.111 0.151 0.104
SWP 0.134 0.110 0.109 0.112 0.073 0.069 0.079 0.167 0.086 0.144 0.094 0.125 0.039 0.145 0.067 0.092 0.153 0.044 0.068 0.103 0.161 0.217
Soybeans CEA 0.193 0.205 0.251 0.235 0.138 0.159 0.140 0.271 0.244 0.257 0.206 0.262 0.205 0.225 0.175 0.224 0.099 0.141 0.097 0.089 0.123 0.155
MID 0.234 0.290 0.378 0.273 0.214 0.268 0.231 0.314 0.288 0.282 0.345 0.286 0.258 0.278 0.248 0.233 0.233 0.243 0.248 0.278 0.243 0.379
SEA 0.193 0.279 0.339 0.262 0.221 0.276 0.234 0.294 0.299 0.305 0.324 0.263 0.235 0.240 0.186 0.374 0.093 0.248 0.106 0.107 0.108 0.238
SPL 0.173 0.221 0.265 0.221 0.211 0.249 0.185 0.169 0.247 0.236 0.230 0.145 0.203 0.184 0.194 0.182 0.077 0.199 0.104 0.108 0.141 0.160
[…]
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.673 0.800 0.684 0.726 0.849 0.782 0.738 0.606 0.785 0.753 0.761 0.805 0.771 0.779 0.836 0.760 0.874 0.864 0.861 0.883 0.891 0.811
CWP 0.795 0.744 0.737 0.817 0.766 0.746 0.717 0.748 0.711 0.835 0.748 0.815 0.854 0.890 0.841 0.844 0.807 0.848 0.841 0.827 0.816 0.728
MID 0.587 0.670 0.709 0.687 0.705 0.741 0.733 0.630 0.732 0.692 0.500 0.807 0.769 0.703 0.700 0.730 0.703 0.717 0.730 0.700 0.717 0.669
SEA 0.749 0.794 0.673 0.758 0.746 0.771 0.785 0.656 0.684 0.665 0.707 0.610 0.781 0.800 0.698 0.761 0.922 0.679 0.899 0.910 0.872 0.816
SPL 0.769 0.677 0.647 0.704 0.617 0.681 0.763 0.627 0.686 0.738 0.647 0.710 0.768 0.732 0.805 0.770 0.941 0.750 0.942 0.954 0.949 0.865
SWP 0.858 0.800 0.697 0.824 0.857 0.875 0.880 0.871 0.794 0.720 0.885 0.904 0.838 0.931 0.800 0.877 0.903 0.880 0.926 0.910 0.891 0.838
Wheat (Winter) CEA 0.735 0.790 0.714 0.829 0.705 0.789 0.830 0.737 0.829 0.571 0.729 0.757 0.742 0.751 0.860 0.774 0.909 0.832 0.887 0.898 0.921 0.845
CWP 0.632 0.804 0.667 0.828 0.762 0.766 0.751 0.718 0.802 0.838 0.732 0.892 0.861 0.893 0.878 0.850 0.841 0.847 0.819 0.824 0.732 0.689
MID 0.706 0.771 0.685 0.689 0.681 0.696 0.721 0.662 0.645 0.692 0.664 0.718 0.698 0.690 0.753 0.721 0.690 0.747 0.753 0.721 0.747 0.745
SEA 0.762 0.695 0.598 0.741 0.710 0.677 0.765 0.727 0.651 0.605 0.743 0.714 0.629 0.661 0.722 0.662 0.906 0.790 0.876 0.865 0.850 0.848
SPL 0.734 0.714 0.612 0.697 0.754 0.683 0.763 0.729 0.723 0.713 0.726 0.730 0.687 0.674 0.661 0.712 0.894 0.683 0.895 0.872 0.853 0.807
SWP 0.779 0.760 0.675 0.731 0.766 0.813 0.827 0.776 0.780 0.755 0.742 0.877 0.827 0.811 0.843 0.872 0.813 0.813 0.824 0.865 0.848 0.796
Soybeans CEA 0.685 0.779 0.686 0.733 0.839 0.834 0.815 0.649 0.784 0.727 0.754 0.803 0.822 0.773 0.793 0.720 0.893 0.828 0.883 0.864 0.876 0.810
MID 0.604 0.707 0.712 0.681 0.745 0.769 0.751 0.664 0.792 0.741 0.563 0.857 0.740 0.754 0.745 0.773 0.754 0.734 0.745 0.773 0.734 0.690
SEA 0.721 0.731 0.726 0.702 0.809 0.770 0.780 0.714 0.648 0.663 0.679 0.574 0.800 0.805 0.756 0.770 0.914 0.651 0.890 0.867 0.864 0.705
SPL 0.692 0.676 0.700 0.641 0.702 0.718 0.790 0.635 0.682 0.797 0.654 0.726 0.783 0.713 0.786 0.811 0.885 0.795 0.927 0.885 0.880 0.820
Table B7: MAPE values of XGB models for each crop and region (Test Set)
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.295 0.251 0.264 0.295 0.166 0.261 0.189 0.326 0.260 0.324 0.267 0.308 0.175 0.301 0.166 0.300 0.119 0.146 0.122 0.194 0.138 0.262
CWP 0.230 0.269 0.351 0.254 0.279 0.247 0.302 0.318 0.282 0.280 0.369 0.253 0.237 0.236 0.191 0.267 0.234 0.219 0.179 0.237 0.272 0.226
MID 0.282 0.310 0.367 0.312 0.222 0.301 0.273 0.352 0.292 0.314 0.370 0.264 0.250 0.345 0.301 0.312 0.345 0.244 0.301 0.312 0.244 0.387
SEA 0.217 0.298 0.419 0.266 0.251 0.369 0.219 0.231 0.269 0.359 0.326 0.284 0.266 0.309 0.217 0.344 0.086 0.314 0.154 0.145 0.111 0.257
SPL 0.173 0.205 0.282 0.194 0.228 0.270 0.201 0.183 0.215 0.312 0.167 0.154 0.182 0.173 0.225 0.202 0.048 0.208 0.052 0.042 0.057 0.067
SWP 0.112 0.108 0.243 0.038 0.081 0.130 0.088 0.043 0.114 0.274 0.036 0.079 0.190 0.092 0.077 0.140 0.035 0.073 0.052 0.035 0.076 0.017
Wheat (Winter) CEA 0.162 0.237 0.329 0.176 0.226 0.249 0.198 0.228 0.211 0.340 0.277 0.257 0.250 0.233 0.171 0.250 0.140 0.179 0.141 0.199 0.141 0.181
CWP 0.179 0.198 0.177 0.177 0.175 0.182 0.168 0.165 0.245 0.098 0.270 0.110 0.163 0.162 0.128 0.199 0.191 0.157 0.135 0.236 0.222 0.208
MID 0.206 0.301 0.365 0.280 0.271 0.306 0.279 0.372 0.344 0.278 0.369 0.316 0.284 0.337 0.302 0.351 0.337 0.277 0.302 0.351 0.277 0.371
SEA 0.249 0.281 0.379 0.292 0.290 0.317 0.193 0.268 0.251 0.321 0.282 0.261 0.286 0.373 0.127 0.300 0.093 0.244 0.180 0.158 0.147 0.170
SPL 0.167 0.236 0.231 0.213 0.161 0.178 0.207 0.193 0.182 0.239 0.206 0.220 0.210 0.215 0.185 0.185 0.065 0.225 0.060 0.068 0.085 0.101
SWP 0.055 0.102 0.132 0.064 0.075 0.070 0.061 0.092 0.053 0.085 0.071 0.082 0.058 0.104 0.050 0.108 0.087 0.070 0.058 0.091 0.094 0.151
Soybeans CEA 0.185 0.230 0.248 0.269 0.156 0.212 0.177 0.286 0.251 0.275 0.226 0.266 0.140 0.264 0.169 0.229 0.123 0.159 0.092 0.142 0.101 0.231
MID 0.237 0.272 0.314 0.271 0.199 0.271 0.236 0.309 0.261 0.306 0.339 0.236 0.229 0.318 0.273 0.257 0.236 0.236 0.273 0.257 0.318 0.367
SEA 0.222 0.314 0.330 0.249 0.239 0.315 0.235 0.283 0.264 0.304 0.303 0.304 0.258 0.323 0.198 0.328 0.112 0.299 0.117 0.088 0.087 0.265
SPL 0.209 0.225 0.263 0.199 0.210 0.276 0.174 0.214 0.257 0.265 0.191 0.192 0.196 0.191 0.218 0.185 0.068 0.217 0.085 0.071 0.088 0.132
[…]
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.206 0.339 0.298 0.307 0.476 0.545 0.416 0.166 0.411 0.335 0.431 0.470 0.497 0.401 0.438 0.427 0.646 0.476 0.619 0.567 0.535 0.316
CWP 0.442 0.343 0.261 0.405 0.291 0.364 0.370 0.387 0.274 0.519 0.409 0.526 0.530 0.557 0.499 0.592 0.421 0.558 0.608 0.556 0.408 0.315
MID 0.151 0.252 0.246 0.255 0.338 0.408 0.388 0.121 0.283 0.318 0.119 0.444 0.249 0.236 0.312 0.268 0.236 0.320 0.312 0.268 0.320 0.205
SEA 0.253 0.239 0.168 0.341 0.285 0.331 0.347 0.113 0.296 0.282 0.158 0.211 0.485 0.379 0.249 0.377 0.658 0.248 0.715 0.671 0.573 0.507
SPL 0.325 0.192 0.192 0.235 0.190 0.236 0.353 0.240 0.337 0.330 0.172 0.336 0.356 0.373 0.402 0.362 0.797 0.366 0.803 0.707 0.825 0.573
SWP 0.582 0.492 0.151 0.571 0.516 0.552 0.622 0.553 0.371 0.250 0.599 0.636 0.617 0.619 0.508 0.549 0.666 0.516 0.641 0.719 0.478 0.520
Wheat (Winter) CEA 0.258 0.489 0.094 0.498 0.335 0.429 0.469 0.354 0.496 0.167 0.308 0.291 0.403 0.354 0.580 0.455 0.618 0.531 0.631 0.709 0.658 0.492
CWP 0.173 0.349 0.201 0.375 0.320 0.285 0.244 0.433 0.349 0.530 0.408 0.629 0.603 0.576 0.547 0.485 0.549 0.533 0.595 0.572 0.392 0.294
MID 0.250 0.302 0.148 0.240 0.190 0.195 0.239 0.206 0.177 0.171 0.172 0.175 0.219 0.175 0.353 0.273 0.175 0.390 0.353 0.273 0.390 0.297
SEA 0.351 0.362 0.135 0.311 0.255 0.258 0.255 0.308 0.278 0.129 0.284 0.236 0.167 0.173 0.323 0.212 0.577 0.216 0.609 0.490 0.488 0.598
SPL 0.284 0.252 0.245 0.241 0.353 0.321 0.298 0.256 0.283 0.341 0.251 0.328 0.363 0.301 0.309 0.335 0.657 0.253 0.666 0.626 0.607 0.518
SWP 0.417 0.365 0.152 0.411 0.345 0.437 0.412 0.408 0.324 0.224 0.308 0.545 0.471 0.380 0.392 0.411 0.570 0.522 0.456 0.560 0.506 0.440
Soybeans CEA 0.226 0.360 0.300 0.323 0.492 0.631 0.494 0.199 0.377 0.311 0.502 0.458 0.516 0.437 0.427 0.435 0.626 0.432 0.630 0.555 0.525 0.370
MID 0.181 0.281 0.276 0.327 0.394 0.478 0.474 0.114 0.386 0.358 0.162 0.504 0.346 0.276 0.320 0.324 0.320 0.371 0.276 0.324 0.371 0.200
SEA 0.155 0.137 0.205 0.247 0.303 0.331 0.332 0.060 0.192 0.287 0.092 0.184 0.465 0.318 0.273 0.383 0.604 0.228 0.534 0.543 0.470 0.302
SPL 0.326 0.229 0.250 0.246 0.239 0.265 0.399 0.209 0.302 0.347 0.183 0.322 0.418 0.385 0.454 0.404 0.684 0.360 0.709 0.617 0.736 0.500
Table B9: MAPE values of LGBM models for each crop and region (Test Set)
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.301 0.273 0.325 0.260 0.270 0.255 0.255 0.357 0.222 0.325 0.270 0.309 0.233 0.311 0.247 0.273 0.220 0.287 0.213 0.249 0.243 0.254
CWP 0.246 0.217 0.332 0.267 0.265 0.260 0.264 0.241 0.276 0.277 0.215 0.238 0.207 0.219 0.254 0.312 0.264 0.286 0.197 0.261 0.334 0.281
MID 0.303 0.320 0.322 0.309 0.253 0.293 0.317 0.330 0.301 0.311 0.336 0.280 0.306 0.332 0.323 0.291 0.332 0.275 0.323 0.291 0.275 0.362
SEA 0.327 0.352 0.422 0.355 0.341 0.365 0.320 0.402 0.362 0.367 0.363 0.358 0.283 0.305 0.342 0.379 0.256 0.360 0.242 0.231 0.253 0.267
SPL 0.152 0.216 0.270 0.209 0.172 0.284 0.211 0.168 0.177 0.272 0.194 0.135 0.196 0.172 0.233 0.217 0.080 0.235 0.062 0.101 0.056 0.087
SWP 0.129 0.155 0.242 0.126 0.139 0.195 0.122 0.049 0.129 0.230 0.054 0.096 0.199 0.086 0.086 0.166 0.060 0.153 0.072 0.045 0.073 0.071
Wheat (Winter) CEA 0.250 0.216 0.332 0.189 0.234 0.242 0.213 0.246 0.235 0.319 0.250 0.268 0.230 0.244 0.221 0.278 0.194 0.234 0.184 0.202 0.213 0.196
CWP 0.182 0.215 0.203 0.179 0.201 0.244 0.182 0.187 0.209 0.138 0.237 0.125 0.161 0.132 0.165 0.193 0.173 0.221 0.135 0.239 0.221 0.183
MID 0.257 0.268 0.300 0.262 0.315 0.287 0.268 0.313 0.302 0.303 0.301 0.306 0.322 0.298 0.291 0.267 0.298 0.247 0.291 0.267 0.247 0.344
SEA 0.238 0.262 0.356 0.276 0.325 0.261 0.296 0.260 0.265 0.308 0.301 0.304 0.288 0.284 0.234 0.285 0.166 0.267 0.173 0.252 0.222 0.219
SPL 0.138 0.207 0.237 0.192 0.195 0.228 0.210 0.160 0.210 0.215 0.191 0.186 0.206 0.200 0.165 0.201 0.096 0.198 0.100 0.094 0.098 0.116
SWP 0.081 0.099 0.066 0.064 0.058 0.053 0.040 0.107 0.075 0.063 0.074 0.048 0.064 0.028 0.030 0.083 0.071 0.079 0.048 0.091 0.110 0.105
Soybeans CEA 0.263 0.240 0.295 0.227 0.228 0.209 0.216 0.305 0.204 0.290 0.231 0.279 0.201 0.272 0.225 0.232 0.215 0.251 0.187 0.213 0.196 0.213
MID 0.259 0.289 0.294 0.262 0.214 0.261 0.266 0.308 0.254 0.292 0.299 0.257 0.279 0.293 0.305 0.244 0.233 0.233 0.305 0.244 0.293 0.336
SEA 0.321 0.323 0.331 0.345 0.286 0.295 0.287 0.385 0.345 0.318 0.342 0.321 0.262 0.300 0.307 0.320 0.247 0.297 0.233 0.241 0.248 0.282
SPL 0.152 0.210 0.252 0.218 0.195 0.255 0.211 0.175 0.194 0.238 0.212 0.167 0.170 0.178 0.193 0.190 0.094 0.196 0.115 0.134 0.130 0.132
[…]
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.654 0.673 0.673 0.769 0.843 0.860 0.878 0.678 0.712 0.789 0.802 0.867 0.858 0.884 0.950 0.883 0.958 0.946 0.959 0.962 0.960 0.777
CWP 0.749 0.708 0.700 0.826 0.847 0.836 0.859 0.817 0.755 0.850 0.860 0.947 0.918 0.957 0.956 0.962 0.956 0.960 0.958 0.962 0.959 0.730
MID 0.682 0.683 0.750 0.792 0.813 0.826 0.852 0.636 0.743 0.742 0.802 0.841 0.802 0.857 0.927 0.908 0.857 0.931 0.927 0.908 0.931 0.618
SEA 0.746 0.727 0.675 0.801 0.835 0.831 0.878 0.653 0.702 0.634 0.747 0.823 0.836 0.847 0.907 0.866 0.975 0.934 0.979 0.976 0.978 0.863
SPL 0.782 0.736 0.617 0.857 0.843 0.811 0.866 0.694 0.682 0.730 0.762 0.875 0.816 0.883 0.913 0.892 0.983 0.920 0.983 0.983 0.984 0.896
SWP 0.835 0.814 0.560 0.887 0.910 0.885 0.928 0.876 0.774 0.758 0.893 0.925 0.886 0.937 0.946 0.946 0.971 0.955 0.970 0.973 0.974 0.797
Wheat (Winter) CEA 0.711 0.825 0.546 0.864 0.825 0.863 0.893 0.758 0.811 0.711 0.850 0.831 0.869 0.893 0.942 0.908 0.961 0.952 0.970 0.963 0.971 0.846
CWP 0.659 0.746 0.747 0.792 0.827 0.884 0.901 0.801 0.779 0.878 0.851 0.938 0.930 0.955 0.964 0.958 0.959 0.966 0.967 0.960 0.967 0.782
MID 0.687 0.702 0.704 0.802 0.746 0.797 0.848 0.696 0.744 0.698 0.800 0.816 0.819 0.853 0.891 0.876 0.853 0.920 0.891 0.876 0.920 0.625
SEA 0.662 0.757 0.655 0.774 0.789 0.848 0.859 0.611 0.687 0.605 0.769 0.777 0.790 0.818 0.914 0.824 0.962 0.919 0.961 0.960 0.966 0.803
SPL 0.763 0.687 0.651 0.800 0.873 0.847 0.887 0.698 0.725 0.744 0.752 0.837 0.846 0.861 0.890 0.883 0.956 0.905 0.961 0.965 0.965 0.832
SWP 0.739 0.760 0.656 0.852 0.874 0.854 0.897 0.747 0.787 0.807 0.846 0.915 0.898 0.929 0.930 0.944 0.966 0.951 0.970 0.967 0.970 0.748
Soybeans CEA 0.674 0.730 0.718 0.814 0.831 0.890 0.896 0.696 0.724 0.783 0.812 0.871 0.879 0.905 0.942 0.901 0.961 0.942 0.949 0.952 0.961 0.797
MID 0.733 0.720 0.738 0.838 0.818 0.865 0.878 0.649 0.798 0.726 0.799 0.863 0.836 0.884 0.927 0.920 0.927 0.946 0.884 0.920 0.946 0.627
SEA 0.691 0.681 0.744 0.783 0.839 0.829 0.867 0.502 0.717 0.710 0.762 0.772 0.818 0.833 0.922 0.858 0.969 0.938 0.966 0.968 0.971 0.819
SPL 0.752 0.681 0.693 0.841 0.848 0.851 0.902 0.671 0.684 0.781 0.731 0.872 0.850 0.893 0.917 0.895 0.969 0.915 0.964 0.972 0.976 0.845
Table B11: MAPE values of XGB-C models for each crop and region (Test Set)
Crop Region Model
1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22
Corn (Grain) CEA 0.247 0.281 0.328 0.280 0.196 0.243 0.206 0.317 0.232 0.373 0.254 0.316 0.249 0.275 0.172 0.285 0.155 0.181 0.150 0.148 0.149 0.237
CWP 0.217 0.222 0.348 0.222 0.257 0.267 0.244 0.250 0.241 0.281 0.225 0.251 0.256 0.235 0.221 0.240 0.234 0.226 0.232 0.245 0.230 0.304
MID 0.271 0.329 0.321 0.302 0.301 0.332 0.329 0.382 0.306 0.373 0.313 0.362 0.334 0.308 0.281 0.310 0.308 0.275 0.281 0.310 0.275 0.398
SEA 0.289 0.269 0.381 0.298 0.293 0.342 0.281 0.347 0.292 0.401 0.291 0.246 0.227 0.192 0.111 0.179 0.051 0.096 0.057 0.056 0.059 0.181
SPL 0.145 0.261 0.283 0.175 0.164 0.218 0.148 0.218 0.232 0.284 0.195 0.154 0.194 0.150 0.224 0.173 0.049 0.225 0.051 0.055 0.055 0.049
SWP 0.027 0.100 0.277 0.048 0.031 0.073 0.042 0.045 0.106 0.246 0.032 0.082 0.088 0.076 0.066 0.077 0.049 0.080 0.043 0.049 0.044 0.050
Wheat (Winter) CEA 0.152 0.172 0.364 0.166 0.170 0.170 0.160 0.268 0.232 0.351 0.244 0.267 0.238 0.236 0.124 0.248 0.156 0.140 0.136 0.145 0.128 0.160
CWP 0.176 0.187 0.201 0.174 0.198 0.186 0.208 0.166 0.162 0.141 0.170 0.152 0.131 0.137 0.113 0.137 0.134 0.120 0.119 0.134 0.117 0.174
MID 0.212 0.317 0.314 0.277 0.274 0.299 0.264 0.375 0.302 0.344 0.269 0.333 0.278 0.270 0.244 0.276 0.270 0.223 0.244 0.276 0.223 0.369
SEA 0.242 0.204 0.350 0.235 0.256 0.236 0.224 0.265 0.238 0.304 0.268 0.203 0.200 0.225 0.121 0.249 0.087 0.112 0.100 0.099 0.101 0.134
SPL 0.143 0.212 0.226 0.180 0.163 0.207 0.173 0.111 0.221 0.251 0.202 0.171 0.225 0.170 0.181 0.178 0.056 0.163 0.063 0.058 0.064 0.088
SWP 0.069 0.088 0.148 0.091 0.078 0.060 0.077 0.067 0.058 0.067 0.094 0.078 0.044 0.086 0.068 0.070 0.087 0.067 0.082 0.088 0.083 0.071
Soybeans CEA 0.199 0.214 0.297 0.217 0.205 0.212 0.187 0.291 0.174 0.331 0.210 0.281 0.225 0.252 0.181 0.251 0.100 0.183 0.100 0.106 0.095 0.199
MID 0.239 0.292 0.297 0.292 0.255 0.302 0.290 0.343 0.299 0.331 0.305 0.321 0.293 0.279 0.275 0.276 0.279 0.249 0.275 0.276 0.249 0.371
SEA 0.326 0.337 0.311 0.392 0.255 0.295 0.280 0.375 0.275 0.308 0.243 0.232 0.239 0.202 0.142 0.195 0.088 0.125 0.091 0.093 0.093 0.228
SPL 0.140 0.259 0.234 0.158 0.172 0.229 0.174 0.166 0.265 0.243 0.185 0.207 0.221 0.201 0.237 0.198 0.068 0.233 0.069 0.086 0.083 0.112